OpenStudy (anonymous):

Carla's plant was 10 2/3 inches tall. Yolanda's was 8 3/4 inches tall. How much taller was Carla's plant than Yolanda's? A. 1 11/12 in. B. 2 1/12 in. C. 2 5/12 in. D. 2 11/12

OpenStudy (anonymous):

i think its c

OpenStudy (superflytyguy):

Did you find the Greatest Common Denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

...

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (superflytyguy):

Okay well 3 and 4 have 12 in common right? Make both denominators the same. 10 2/3 = 10 8/12 8 3/4 = 8 9/12 Subtract 10 8/12 - 8 9/12 10 8/12 = 128/12 8 9/12 = 105/12 128 - 105 = 23 23/12 1 11/12

OpenStudy (anonymous):

okay look

OpenStudy (superflytyguy):

A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):